# Heterogeneous Directed Hypergraph Neural Network over abstract syntax tree (AST) for Code Classification

### v4

@@ Abstract
Code classification is a difficult issue in program understanding and automatic coding. Due to the elusive syntax and complicated semantics in programs, most existing studies use techniques based on abstract syntax tree (AST) and graph neural networks (GNN) to create code representations for code classification. These techniques utilize the structure and semantic information of the code, but they only take into account pairwise associations and neglect the high-order data correlations that already exist between nodes of the same field or called attribute in the AST, which may result in the loss of code structural information. On the other hand, while a general hypergraph can encode high-order data correlations, it is homogeneous and undirected which will result in a lack of semantic and structural information such as node types, edge types, and directions between child nodes and parent nodes when modeling AST. In this study, we propose a heterogeneous directed hypergraph (HDHG) to represent AST and a heterogeneous directed hypergraph neural network (HDHGN) to process the graph for code classification. Our method improves code understanding and can represent high-order data correlations beyond paired interactions. We assess our heterogeneous directed hypergraph neural network (HDHGN) on public datasets of Python and Java programs. Our method outperforms previous AST-based and GNN-based methods, which demonstrates the capability of our model.
Index Terms:
hypergraph, heterogeneous graph, code classification, graph neural networks, code representation
@@ I Introduction
Due to the improvement of machine learning technology, particularly deep learning, more and more work has employed deep learning for code classification. Currently, there are two main categories of code classification methods: AST-based and GNN-based. To take advantage of the semantic and structural information of the source code, several studies adopt AST when learning code representations [1][5][7][8]. Some research uses graph neural networks (GNN) to create code representations for code categorization to a better understanding of the structure of code based on AST [3][4][13][14].
@@ II Related Work
With the popularity of GNN, more works apply kinds of GNN in code classification based on AST to strengthen the comprehension of code structures. M. Allamanis et al. [13] first construct graphs from source code by adding edges like control flow and data flow to AST and employing a gated graph neural network (GGNN) to process program graphs. V. Hellendoorn et al. [19] propose a model called GREAT based on the transformer architecture by extracting global relational information from code graphs. D. Vagavolu et al. [3] propose an approach that can extract and use program features from multiple code graphs. M. Lu et al. [6] improve gated graph neural network (GGNN) in program classification. T. Long et al. [14] propose a multi-view graph program representation method that combines both data flow and control flow as multiple views and applies GNN to process. W. Wang et al. [4] propose to leverage heterogeneous graphs to show code based on previous work and adopt heterogeneous GNN to process.
@@ IV-B1 Definition
We let a HDHG $G=(N,E)$, which includes a node set $N=\left\{n_{1},~n_{2},\ldots,n_{|N|}\right\}$ and a directed hyperedge set $E=\left\{e_{1},e_{2},\ldots,e_{|E|}\right\}$. Each node $n=(\mu,x)$, where $\mu$ represents node type and $x$ is the value of the node. Each directed hyperedge $e=\left({\rho,S(e),T(e)}\right)$, $\rho$ represents edge type, $S(e)=\left\{n_{1},\ldots,n_{|{S(e)}|}\right\}\subseteq N$ is the tail nodes of hyperedge $e$, $T(e)\in N$ is the head node of hyperedge $e$, they show the direction of the hyperedge $e$ is $S(e)$ to $T(e)$.
@@ V-D Results
HPG+HGT444 We use the outcomes that were reported in their research because the paper did not make their code publicly available. 94.99 - 93.95 -

### v6

@@ Abstract
Code classification is a difficult issue in program understanding and automatic coding. Due to the elusive syntax and complicated semantics in programs, most existing studies use techniques based on abstract syntax tree (AST) and graph neural network (GNN) to create code representations for code classification. These techniques utilize the structure and semantic information of the code, but they only take into account pairwise associations and neglect the high-order data correlations that already exist between nodes of the same field or called attribute in the AST, which may result in the loss of code structural information. On the other hand, while a general hypergraph can encode high-order data correlations, it is homogeneous and undirected which will result in a lack of semantic and structural information such as node types, edge types, and directions between child nodes and parent nodes when modeling AST. In this study, we propose a heterogeneous directed hypergraph (HDHG) to represent AST and a heterogeneous directed hypergraph neural network (HDHGN) to process the graph for code classification. Our method improves code understanding and can represent high-order data correlations beyond paired interactions. We assess our heterogeneous directed hypergraph neural network (HDHGN) on public datasets of Python and Java programs. Our method outperforms previous AST-based and GNN-based methods, which demonstrates the capability of our model.
@@ I Introduction
Due to the improvement of machine learning technology, particularly deep learning, more and more work has employed deep learning for code classification. Currently, there are two main categories of code classification methods: AST-based and GNN-based. To take advantage of the semantic and structural information of the source code, several studies adopt AST when learning code representations [1][5][7][8]. Some research uses graph neural network (GNN) to create code representations for code categorization to a better understanding of the structure of code based on AST [3][4][13][14].
@@ II Related Work
With the popularity of GNN, more works apply kinds of GNN in code classification based on AST to strengthen the comprehension of code structures. M. Allamanis et al. [13] first construct graphs from source code by adding edges like control flow and data flow to AST and employing a gated graph neural network (GGNN) to process program graphs. V. Hellendoorn et al. [19] propose a model called GREAT based on the transformer architecture by extracting global relational information from code graphs. D. Vagavolu et al. [3] propose an approach that can extract and use program features from multiple code graphs. M. Lu et al. [6] improve gated graph neural network (GGNN) in program classification. T. Long et al. [14] propose a multi-view graph program representation method that combines both data flow and control flow as multiple views and apply GNN to process. W. Wang et al. [4] propose to leverage heterogeneous graphs to show code based on previous work and adopt heterogeneous GNN to process.
@@ IV-B1 Definition
We let an HDHG $G=(N,E)$, which includes a node set $N=\left\{n_{1},~n_{2},\ldots,n_{|N|}\right\}$ and a directed hyperedge set $E=\left\{e_{1},e_{2},\ldots,e_{|E|}\right\}$. Each node $n=(\mu,x)$, where $\mu$ represents node type and $x$ is the value of the node. Each directed hyperedge $e=\left({\rho,S(e),T(e)}\right)$, $\rho$ represents edge type, $S(e)=\left\{n_{1},\ldots,n_{|{S(e)}|}\right\}\subseteq N$ is the tail nodes of hyperedge $e$, $T(e)\in N$ is the head node of hyperedge $e$, they show the direction of the hyperedge $e$ is $S(e)$ to $T(e)$.
@@ V-D Results
HPG+HGT [^4] We use the outcomes that were reported in their research because the paper did not make their code publicly available. 94.99 - 93.95 -